The probability of going to a resort and winning a new car is 1 in 52,000.
If 5,000 people visit a resort, how many would be expected to win a new car?

Answer by stanbon(75887)

The probability of going to a resort and winning a new car is 1 in 52,000.
If 5,000 people visit a resort, how many would be expected to win a new car?
-------
E(x) = n*p = 5000(1/52000) = 0.0962 person
------------
Cheers,
Stan H.